Problem 01.034 Simple and Compound Interest Durco Automotive needs a $1 million balance in its contingency fund 3 years from now. The chief financial officer (CFO) wants to know how much to deposit now into Durco’s high-yield investment account. Problem 01.034.b Simple and Compound Interest Determine the amount if it grows at a rate of 11% per year using compound interest. The amount if it grows at a rate of 11% per year using compound interest is $
Answer #1
A B C D E F G H I J K L
2
3 Using compound interest rate i, the future value after year n of amount A deposited today can be calculated as follows:
4 FV =A*(1+i)n
5
6 Given the following,
7 FV $1,000,000
8 i 11%
9 n 3
10
11 A =FV/ (1+i)n
12 $731,191.38 =D7/((1+D8)^D9)
13
14 Hence Amount to be deposited today is $731,191.38
15